# Fast post-process Bayesian inference with Variational Sparse Bayesian Quadrature

### v2

@@ 1 Introduction
Due to the cost of inference, the workflow of Bayesian analyses often starts with a preliminary exploration phase via simpler and cheaper means (Gelman et al., 2020). A popular choice consists of performing maximum a posteriori (MAP) estimation, i.e., finding the (global) mode of the posterior density, often via multiple runs of black-box optimization algorithms (e.g., Hansen et al., 2003; Acerbi and Ma, 2017), which can easily require tens of thousands likelihood or posterior density evaluations over distinct optimization runs to satisfactorily explore the parameter landscape (e.g., Acerbi et al., 2018; Norton et al., 2019; Cao et al., 2019; Zhou et al., 2020; Yoo et al., 2021; Heald et al., 2021). In fact, due to prohibitive costs, many analyses stop here, with a point estimate instead of a full posterior, as advocated for example in modeling tutorials and textbooks in applied domains such as computational and cognitive neuroscience (Wilson and Collins, 2019; Ma et al., 2023).
@@
As an instantiation of our framework, we introduce Variational Sparse Bayesian Quadrature (vsbq). vsbq builds a Gaussian process (GP; Rasmussen and Williams, 2006) surrogate of the log density, starting from existing log-likelihood or log-density evaluations. In particular, we use a sparse GP to deal with a potentially large number of model evaluations (Titsias, 2009). A tractable posterior approximation is then obtained by performing variational inference over the surrogate, thus without further evaluations of the original model. Fitting the flexible variational posterior is particularly efficient by utilizing Bayesian quadrature (O’Hagan, 1991; Acerbi, 2018). We validate our approach on synthetic targets and real datasets and models from computational neuroscience (Acerbi et al., 2012, 2018). Overall, we find that post-process Bayesian inference via vsbq is not only feasible but can yield high-quality approximations, providing the applied modeler with a new approximate inference tool that can easily fit in existing modeling pipelines as a refinement step. At the end of the paper, we discuss the limitations of the method and future work.
@@
There are several works related to post-processing existing density evaluations to construct an approximate posterior. Zhang et al. (2022) find the best multivariate normal approximation along the optimization paths generated by a quasi-Newton optimization algorithm, L-BFGS (Liu and Nocedal, 1989), in terms of Kullback-Leibler (KL) divergence to the true posterior. Unlike our method, the L-BFGS algorithm requires the gradients of the log-likelihood or log-density function to build approximate estimates of the Hessian along the optimization trajectory, and is brittle to noise in the target function. Moreover, the post-process estimation of the KL divergence needs additional evaluations of the log-density function. Yao et al. (2022) propose to reuse parallel – and possibly incomplete – runs of different inference algorithms such as Markov Chain Monte Carlo (mcmc; Robert and Casella, 2004, see e.g.) or variational inference (Blei et al., 2017) by combining them in a weighted average via ‘Bayesian stacking’. Similar in spirit to our approach, the computation of the stacking weights only requires a post-processing step. However, unlike our method, Bayesian stacking requires the samples to be approximately drawn from the posterior and does not make use of all available target evaluations. A related technique that merits mention is offline black-box optimization (Krishnamoorthy et al., 2023; Trabucco et al., 2022). In offline black-box optimization, the objective is to identify a parameter input that maximizes a black-box function utilizing pre-existing offline function evaluations. Much like our approach, offline black-box optimization capitalizes on leveraging existing evaluations. However, a key distinction lies in our primary objective, which is to construct an approximation of the unknown posterior density, a task inherently more challenging than finding a point estimate.
@@ 2.1 Variational inference
where $\mathbb{E}_{\bm{\phi}}\equiv\mathbb{E}_{q_{\bm{\phi}}}$. Crucially, $D_{\text{KL}}(q||p)\geq 0$ and $D_{\text{KL}}(q||p)=0$ if and only if $q=p$. Minimizing the KL divergence casts Bayesian inference as an optimization problem. This optimization is finding the variational parameters, $\bm{\phi}$, that maximizes the objective:
$\text{{elbo}}\left[q_{\bm{\phi}}\right]=\mathbb{E}_{\bm{\phi}}\left[\log\frac{p(\mathcal{D}|\mathbf{x})p(\mathbf{x})}{q_{\bm{\phi}}(\mathbf{x})}\right]=\mathbb{E}_{\bm{\phi}}\left[f_{0}(\mathbf{x})\right]+\mathcal{H}[q_{\bm{\phi}}(\mathbf{x})],$ (2)
@@ 2.2 Gaussian processes
Figure 1: Log-density function modeling and posterior approximation. (a) The left panel depicts the ground-truth (unnormalized) posterior log-density function. The right panel illustrates the Gaussian Process (GP) approximation to the log-density function (mean and 95% credible interval), given observed log-density evaluations (blue crosses). (b) The left panel depicts the ground-truth posterior density function, corresponding to the log density in (a). The right panel illustrates the posterior density approximation corresponding to the GP surrogate.
Covariance function.
In this paper, we consider the exponentiated quadratic kernel,
$\kappa(\mathbf{x},\mathbf{x}^{\prime};\sigma_{f}^{2},\bm{\ell})=\sigma_{f}^{2}\exp\left[-\frac{1}{2}\left(\mathbf{x}-\mathbf{x}^{\prime}\right)\mathbf{\Sigma}_{\ell}^{-1}\left(\mathbf{x}-\mathbf{x}^{\prime}\right)^{\top}\right],$ (3)
@@
The GP posterior given function observations $(\mathbf{X},\mathbf{y},\mathbf{s})$ is also a Gaussian process with posterior mean function $\mu_{\text{p}}$ and posterior covariance function $\kappa_{\text{p}}$ (Rasmussen and Williams, 2006):
$\mu_{\text{p}}(\tilde{\mathbf{x}})$ $=\kappa(\tilde{\mathbf{x}},\mathbf{X})\left(\kappa(\mathbf{X},\mathbf{X})+\mathbf{S}\right)^{-1}\left(\mathbf{y}-m(\mathbf{X})\right)+m(\tilde{\mathbf{x}}),$ (4)
$\kappa_{\text{p}}(\tilde{\mathbf{x}},\tilde{\mathbf{x}}^{\prime})$ $=\kappa(\tilde{\mathbf{x}},\tilde{\mathbf{x}}^{\prime})-\kappa(\tilde{\mathbf{x}},\mathbf{X})\left(\kappa(\mathbf{X},\mathbf{X})+\mathbf{S}\right)^{-1}\kappa(\mathbf{X},\tilde{\mathbf{x}}^{\prime}).$ (5)
@@
A key observation is that a GP surrogate of the log density (Figure 1a) does not immediately yield a usable approximate posterior. The normalization constant is unknown and we cannot directly sample from the density associated with the GP surrogate. Another layer of approximation is needed to go from the log density to the density (Figure 1b). Several approximation techniques have been developed to tackle this issue, many of them involving Bayesian quadrature (Ghahramani and Rasmussen, 2002b; Osborne et al., 2012; Gunter et al., 2014; Acerbi, 2018; Adachi et al., 2022).
@@ 2.3 Bayesian quadrature
where $\mu_{\text{p}}$ and $\kappa_{\text{p}}$ are the GP posterior mean and covariance function. Importantly, if $f$ has a Gaussian kernel and $\pi$ is a Gaussian or mixture of Gaussians (among other functional forms), the integrals in Eq. 7 have closed-form solutions.
@@ 2.4 Variational Bayesian Monte Carlo
Variational Bayesian Monte Carlo (vbmc; Acerbi, 2018, 2020; Huggins et al., 2023) is a framework that combines variational inference (Section 2.1), Gaussian processes (Section 2.2), and Bayesian quadrature (Section 2.3) for approximating the posterior density. vbmc employs a Gaussian process as a surrogate to the log-density function and then performs variational inference on the GP surrogate as opposed to using the original expensive target. The surrogate elbo in variational inference is efficiently estimated via Bayesian quadrature. Moreover, vbmc introduces acquisition functions for actively sampling new evaluations of the log density to iteratively refine the posterior approximation. In this section, we focus on three key components of vbmc related to our work: the surrogate ELBO, the choice of GP mean function, and the specification of the variational posterior.
@@
GP mean function.
[…]
$m(\mathbf{x};m_{0},\bm{\mu},\bm{\omega})\equiv m_{0}-\frac{1}{2}\sum_{i=1}^{D}\frac{\left(x_{i}-\mu_{i}\right)^{2}}{\omega_{i}^{2}},$ (10)
where $m_{0}$ denotes the maximum, $\bm{\mu}\equiv\left(\mu_{1},\ldots,\mu_{D}\right)$ is the location vector, and $\bm{\omega}\equiv\left(\omega_{1},\ldots,\omega_{D}\right)$ is a vector of scale parameters (Acerbi, 2018).
@@
Finally, vbmc takes the variational posterior family to be a flexible mixture of $K$ multivariate Gaussians,
$q_{\bm{\phi}}(\mathbf{x})=\sum_{k=1}^{K}w_{k}\mathcal{N}\left({\mathbf{x}};{\bm{\mu}_{k}},{\sigma_{k}^{2}\mathbf{\Sigma}_{\bm{\lambda}}}\right),\qquad\mathbf{\Sigma}_{\bm{\lambda}}\equiv\text{diag}[{{\lambda}^{(1)}}^{2},\ldots,{\lambda^{(D)}}^{2}],$ (11)
where $w_{k}$, $\bm{\mu}_{k}$, and $\sigma_{k}$ are, respectively, the mixture weight, mean, and global scale of the $k$-th component; and $\bm{\lambda}=({\lambda}^{(1)},\ldots,\lambda^{(D)})$ is a vector of parameter scales shared across mixture components. When coupled with a Gaussian process with the exponentiated quadratic kernel and a negative quadratic mean function, this choice of variational posterior enables closed-form computation of the expected log joint in Eq. 9 via Bayesian quadrature. Consequently, it affords efficient and robust optimization of the variational objective.
@@ 2.5 Sparse Variational Gaussian Processes
For the case of sparse GP regression with Gaussian likelihood, Titsias (2009) proved that the optimal variational distribution that maximizes Eq. 14 for fixed GP hyperparameters and inducing point locations $\mathbf{Z}$ is given by $\tilde{p}(\mathbf{u})=\mathcal{N}(\mathbf{m}_{\mathbf{u}},\mathbf{R}_{\mathbf{u}\mathbf{u}})$, defined below. Furthermore, equality between the sparse and full GP posterior is obtained for $\mathbf{Z}=\mathbf{X}$. We denote with $\bm{\psi}$ the optimal variational parameters $\mathbf{m}_{\mathbf{u}},\mathbf{R}_{\mathbf{u}\mathbf{u}}$ of the sparse GP, defined as:
@@
SGPR posterior.
The sgpr posterior given function observations $(\mathbf{X},\mathbf{y},\mathbf{s})$ and inducing point locations $\mathbf{Z}$ is also a Gaussian process with mean and covariance:
$\kappa_{\bm{\psi}}(\tilde{\mathbf{x}},\tilde{\mathbf{x}}^{\prime})$ $=\kappa(\tilde{\mathbf{x}},\tilde{\mathbf{x}}^{\prime})-\kappa(\tilde{\mathbf{x}},\mathbf{Z})(\mathbf{K}_{\mathbf{Z},\mathbf{Z}}^{-1}-\bm{\Sigma})\kappa(\mathbf{Z},\tilde{\mathbf{x}}^{\prime}).$ (18)
SGPR ELBO.
[…]
where $\mathbf{Q}_{\mathbf{X},\mathbf{X}}\equiv\mathbf{K}_{\mathbf{X},\mathbf{Z}}\mathbf{K}_{\mathbf{Z},\mathbf{Z}}^{-1}\mathbf{K}_{\mathbf{Z},\mathbf{X}}$. Note that the GP-elbo depends on the location of inducing points $\mathbf{Z}$ and GP hyperparameters $\bm{\xi}$, and consists of the log probability density function of a multivariate normal term, minus the trace of an error term. Detailed derivations and numerical implementation details are provided in Supplementary Material A.1-A.3.
@@ 3 Variational Sparse Bayesian Quadrature
In this section, we present our method for post-process Bayesian inference, named Variational Sparse Bayesian Quadrature (vsbq). As mentioned, statistical analysis in computational modeling studies often relies on maximum a posteriori (MAP) estimation, [^2] Often with a uniform, noninformative prior, effectively reducing to maximum-likelihood estimation. typically involving multiple runs of a numerical optimization algorithm to identify the MAP estimate from the highest log-density value. Crucially, the evaluation traces from MAP optimization, which hold valuable information about the posterior density, are usually discarded. The post-process inference framework we propose aims to reuse these many existing evaluations to efficiently construct a good approximation of the posterior distribution, which is particularly beneficial in scenarios with computationally expensive or noisy model evaluations. Our approach recycles valuable information, effectively converting a point estimate into a posterior approximation with minimal computational expense. [^3] An implementation of our algorithm and code used to run our experiments are available at https://github.com/acerbilab/vsbq.
@@ 3.1 Overview of the algorithm
The sparse GP surrogate model uses an exponentiated quadratic kernel (Eq. 3) and a negative quadratic mean function (Eq. 10) to ensure compatibility with Bayesian quadrature and integrability of the exponentiated surrogate. Fitting a sparse GP to the log-density observations involves the selection of inducing points and the sparse GP hyperparameters. We select hyperparameters by maximizing the GP-elbo in Eq. 19. Inducing points selection is critical for the performance of a sparse GP and our chosen approach is detailed in Section 3.2.
@@
As per the vbmc method described in Section 2.4, we take the variational posterior $q_{\phi}$ to be a mixture of $K$ multivariate Gaussians of the form $q_{\bm{\phi}}(\mathbf{x})=\sum_{k=1}^{K}w_{k}\mathcal{N}\left({\mathbf{x}};{\bm{\mu}_{k}},{\sigma_{k}^{2}\mathbf{\Sigma}_{\bm{\lambda}}}\right)$, where $w_{k}$, $\bm{\mu}_{k}$, and $\sigma_{k}$ are the mixture weight, mean, and global scale of the $k$-th component; and $\bm{\Sigma_{\lambda}}$ the common diagonal covariance matrix with a shared vector $\bm{\lambda}$ which specifies a separate scale for each parameter.
@@
$\begin{split}\mathrm{Cov}(\mathcal{I}_{i},\mathcal{I}_{j})=&\int\int\mathcal{N}(\tilde{\mathbf{x}};\bm{\mu}_{i},\bm{\Sigma}_{i})\mathcal{N}(\tilde{\mathbf{x}}^{\prime};\bm{\mu}_{j},\bm{\Sigma}_{j})\mathrm{Cov}(f(\tilde{\mathbf{x}}),f(\tilde{\mathbf{x}}^{\prime}))\mathrm{d}\tilde{\mathbf{x}}\mathrm{d}\tilde{\mathbf{x}}^{\prime}\\ =&\int\int\mathcal{N}(\tilde{\mathbf{x}};\bm{\mu}_{i},\bm{\Sigma}_{i})\mathcal{N}(\tilde{\mathbf{x}}^{\prime};\bm{\mu}_{j},\bm{\Sigma}_{j})\kappa_{\bm{\psi}}(\tilde{\mathbf{x}},\tilde{\mathbf{x}}^{\prime})\mathrm{d}\tilde{\mathbf{x}}\mathrm{d}\tilde{\mathbf{x}}^{\prime},\end{split}$ (23)
where $\kappa_{\bm{\psi}}$ is the sparse GP posterior covariance function as in Eq. 18. Both the posterior mean and variance of the integral can be obtained in closed form as integrals for the product of Gaussians. Therefore, we can compute the elbo and its standard deviation $\textsc{elbo}_{\textrm{sd}}$ efficiently. Thus, once the sparse GP fitting is done, obtaining a tractable posterior approximation is computationally cheap. Derivations are provided in Supplementary Material A.4.

### v3

@@ 1 Introduction
Due to the cost of inference, parameter estimation and the workflow of Bayesian analyses often start with a preliminary exploration phase via simpler and cheaper means (Gelman et al., 2020). A popular choice consists of performing maximum likelihood estimation (MLE) or maximum a posteriori (MAP) estimation (Gelman et al., 2013, Chapter 13), i.e., finding the (global) mode of the posterior density, [^1] For a fixed parameterization, MLE can be viewed as MAP estimation with (improper) uniform/flat priors. often via multiple runs of black-box optimization algorithms (e.g., Hansen et al., 2003; Acerbi and Ma, 2017), which can easily require tens of thousands likelihood or posterior density evaluations over distinct optimization runs to satisfactorily explore the parameter landscape (e.g., Acerbi et al., 2018; Norton et al., 2019; Cao et al., 2019; Zhou et al., 2020; Yoo et al., 2021; Heald et al., 2021). In fact, due to prohibitive costs, many analyses stop here, with a point estimate instead of a full posterior, as advocated for example in modeling tutorials and textbooks in applied domains such as computational and cognitive neuroscience (Wilson and Collins, 2019; Ma et al., 2023).
@@
As an instantiation of our framework, we introduce Variational Sparse Bayesian Quadrature (vsbq). vsbq builds a Gaussian process (GP; Rasmussen and Williams, 2006) surrogate of the log density, starting from existing log-likelihood or log-density evaluations. In particular, we use a sparse GP to deal with a potentially large number of model evaluations (Titsias, 2009) and develop the noise shaping technique for efficient posterior modeling. A tractable posterior approximation is then obtained by performing variational inference over the surrogate, thus without further evaluations of the original model. Fitting the flexible variational posterior is particularly efficient by utilizing Bayesian quadrature (O’Hagan, 1991; Acerbi, 2018). We validate our approach on synthetic targets and real datasets and models from computational neuroscience (Acerbi et al., 2012, 2018). Overall, we find that post-process Bayesian inference via vsbq is not only feasible but can yield high-quality approximations, providing the applied modeler with a new approximate inference tool that can easily fit in existing modeling pipelines as a refinement step. At the end of the paper, we discuss the limitations of the method and future work.
@@
There are several works related to post-processing existing density evaluations to construct an approximate posterior. Zhang et al. (2022) find the best multivariate normal approximation along the optimization paths generated by a quasi-Newton optimization algorithm, L-BFGS (Liu and Nocedal, 1989), in terms of Kullback-Leibler (KL) divergence to the true posterior. Unlike our method, the L-BFGS algorithm requires the gradients of the log-likelihood or log-density function to build approximate estimates of the Hessian along the optimization trajectory and is brittle to noise in the target function. Moreover, the post-process estimation of the KL divergence needs additional evaluations of the log-density function. Yao et al. (2022) propose to reuse parallel – and possibly incomplete – runs of different inference algorithms such as Markov Chain Monte Carlo (mcmc; Robert and Casella, 2004, see e.g.) or variational inference (Blei et al., 2017) by combining them in a weighted average via ‘Bayesian stacking’. Similar in spirit to our approach, the computation of the stacking weights only requires a post-processing step. However, Bayesian stacking requires the samples to be approximately drawn from the posterior and does not make use of all available target evaluations. Finally, a related technique that merits mention is offline black-box optimization (Krishnamoorthy et al., 2023; Trabucco et al., 2022). In offline black-box optimization, the objective is to identify a parameter input that maximizes a black-box function utilizing pre-existing offline function evaluations. Much like our approach, offline black-box optimization capitalizes on leveraging existing evaluations. However, a key distinction lies in our primary objective, which is to construct an approximation of the unknown posterior density, a task inherently more challenging than finding a point estimate.
@@ 2 Background
Figure 1: Log-density function modeling and posterior approximation. (a) The left panel depicts the ground-truth (unnormalized) posterior log-density function. The right panel illustrates the Gaussian Process (GP) approximation to the log-density function (mean and 95% credible interval), given observed log-density evaluations (blue crosses). (b) The left panel depicts the ground-truth posterior density function, corresponding to the log density in (a). The right panel illustrates the posterior density approximation corresponding to the GP surrogate mean.
@@ 2.1 Variational inference
where $\mathbb{E}_{\bm{\phi}}\equiv\mathbb{E}_{q_{\bm{\phi}}}$. Crucially, $D_{\text{KL}}(q||p)\geq 0$ and $D_{\text{KL}}(q||p)=0$ if and only if $q=p$. Minimizing the KL divergence casts Bayesian inference as an optimization problem. This optimization consists of finding the variational parameters, $\bm{\phi}$, that maximize the objective:
$\begin{split}\text{{elbo}}\left[q_{\bm{\phi}}\right]&=\mathbb{E}_{\bm{\phi}}\left[\log\frac{p(\mathcal{D}|\mathbf{x})p(\mathbf{x})}{q_{\bm{\phi}}(\mathbf{x})}\right]\\ &=\mathbb{E}_{\bm{\phi}}\left[f_{0}(\mathbf{x})\right]+\mathcal{H}[q_{\bm{\phi}}(\mathbf{x})],\end{split}$ (2)
@@ 2.2 Gaussian processes
The GP posterior given function observations $(\mathbf{X},\mathbf{y},\mathbf{s})$ is also a Gaussian process with the posterior mean function $\mu_{\text{p}}$ and posterior covariance function $\kappa_{\text{p}}$ (Rasmussen and Williams, 2006). For $\tilde{\mathbf{x}},\tilde{\mathbf{x}}^{\prime}\in\mathcal{X}$,
$\mu_{\text{p}}(\tilde{\mathbf{x}})$ $=\kappa(\tilde{\mathbf{x}},\mathbf{X})\left(\kappa(\mathbf{X},\mathbf{X})+\mathbf{S}\right)^{-1}\left(\mathbf{y}-m(\mathbf{X})\right)$
$\quad+m(\tilde{\mathbf{x}}),$ (3)
$\kappa_{\text{p}}(\tilde{\mathbf{x}},\tilde{\mathbf{x}}^{\prime})$ $=\kappa(\tilde{\mathbf{x}},\tilde{\mathbf{x}}^{\prime})-\kappa(\tilde{\mathbf{x}},\mathbf{X})\left(\kappa(\mathbf{X},\mathbf{X})+\mathbf{S}\right)^{-1}$
×$\quad\times\kappa(\mathbf{X},\tilde{\mathbf{x}}^{\prime}).$ (4)
@@
A key observation is that a GP surrogate of the log density (Figure 1a) does not immediately yield a usable approximate posterior. The normalization constant is unknown and we cannot directly sample from the density associated with the GP surrogate. Another layer of approximation is needed to go from the log density surrogate to posterior approximation (Figure 1). Several approximation techniques have been developed to tackle this issue, many of them involving Bayesian quadrature (Ghahramani and Rasmussen, 2002b; Osborne et al., 2012; Gunter et al., 2014; Acerbi, 2018; Adachi et al., 2022).
@@ 2.3 Bayesian quadrature
[…]
where $\mu_{\text{p}}$ and $\kappa_{\text{p}}$ are the GP posterior mean and covariance function. Importantly, if $f$ has a Gaussian kernel and $\pi$ is a Gaussian or mixture of Gaussians (among other functional forms), the integrals in Eqs. 6 and 7 have closed-form solutions.
@@ 2.4 Variational Bayesian Monte Carlo
Variational Bayesian Monte Carlo (vbmc; Acerbi, 2018, 2020; Huggins et al., 2023) is a framework that combines variational inference (Section 2.1), Gaussian processes (Section 2.2), and Bayesian quadrature (Section 2.3) with the goal of approximating the posterior density. vbmc employs a Gaussian process as a surrogate to the log-density function and then performs variational inference on the GP surrogate as opposed to using the original expensive target. The surrogate elbo in variational inference is efficiently estimated via Bayesian quadrature. Moreover, vbmc introduces acquisition functions for actively sampling new evaluations of the log density to iteratively refine the posterior approximation.
@@
vbmc takes the variational posterior family to be a flexible mixture of multivariate Gaussians. When coupled with a Gaussian process with the exponentiated quadratic kernel and a negative quadratic mean function (see Section 3.1 for the detailed form), this choice of variational posterior enables closed-form computation of the expected log joint in Eq. 9 via Bayesian quadrature. Consequently, it affords efficient and robust optimization of the variational objective. [^3] Note that the entropy of a Gaussian mixture does not admit a closed form, requiring the surrogate elbo to be optimized using stochastic gradient descent.
@@ 2.5 Sparse Variational Gaussian Processes
For the case of sparse GP regression with Gaussian likelihood, Titsias (2009) proved that the optimal variational distribution that maximizes Eq. 12 for fixed GP hyperparameters and inducing point locations $\mathbf{Z}$ is given by $\tilde{p}(\mathbf{u})=\mathcal{N}(\mathbf{m}_{\mathbf{u}},\mathbf{R}_{\mathbf{u}\mathbf{u}})$, defined below. Furthermore, equality between the sparse and full GP posterior is obtained for $\mathbf{Z}=\mathbf{X}$. The optimal variational parameters $\mathbf{m}_{\mathbf{u}},\mathbf{R}_{\mathbf{u}\mathbf{u}}$ of the sparse GP are:
@@
Detailed derivations and numerical implementation details are provided in Supplementary Material A.1-A.3.
@@ 3 Variational Sparse Bayesian Quadrature
In this section, we present our method for post-process Bayesian inference, named Variational Sparse Bayesian Quadrature (vsbq). As mentioned, statistical analysis in computational modeling studies often relies on maximum a posteriori (MAP) estimation, [^4] Often with a uniform, noninformative prior, effectively reducing to maximum-likelihood estimation. typically involving multiple runs of a numerical optimization algorithm to identify the MAP estimate from the highest log-density value. Crucially, the evaluation traces from MAP optimization, which hold valuable information about the posterior density, are usually discarded. The post-process inference framework we propose aims to reuse these many existing evaluations to efficiently construct a good approximation of the posterior distribution, which is particularly beneficial in scenarios with computationally expensive or noisy model evaluations. Our approach recycles valuable information, effectively converting a point estimate into a posterior approximation with minimal computational expense.
@@ 3.1 Overview of the algorithm
The sparse GP surrogate model uses an exponentiated quadratic kernel,
$\begin{split}&\kappa(\mathbf{x},\mathbf{x}^{\prime};\sigma_{f}^{2},\bm{\ell})\\ &=\sigma_{f}^{2}\exp\left[-\frac{1}{2}\left(\mathbf{x}-\mathbf{x}^{\prime}\right)\mathbf{\Sigma}_{\ell}^{-1}\left(\mathbf{x}-\mathbf{x}^{\prime}\right)^{\top}\right],\end{split}$ (16)
The mean function is chosen to be a negative quadratic mean function (Eq. 17) to ensure compatibility with Bayesian quadrature and integrability of the exponentiated surrogate, same as Acerbi (2018),
$m(\mathbf{x};m_{0},\bm{\mu},\bm{\omega})=m_{0}-\frac{1}{2}\sum_{i=1}^{D}\frac{\left(x_{i}-\mu_{i}\right)^{2}}{\omega_{i}^{2}},$ (17)
[…]
The sparse GP posterior given function observations $(\mathbf{X},\mathbf{y},\mathbf{s})$ and inducing point locations $\mathbf{Z}$ is also a Gaussian process with mean and covariance:
$\kappa_{\bm{\psi}}(\tilde{\mathbf{x}},\tilde{\mathbf{x}}^{\prime})$ $=\kappa(\tilde{\mathbf{x}},\tilde{\mathbf{x}}^{\prime})-\kappa(\tilde{\mathbf{x}},\mathbf{Z})(\mathbf{K}_{\mathbf{Z},\mathbf{Z}}^{-1}-\bm{\Sigma})\kappa(\mathbf{Z},\tilde{\mathbf{x}}^{\prime}),$ (19)
where $\bm{\Sigma}$ is defined in Eq. 14.
Fitting a sparse GP to the log-density observations involves two critical components: the selection of inducing points and the sparse GP hyperparameters. Our chosen approach is detailed in Section 3.2.
@@
As per the vbmc method described in Section 2.4, we take the variational posterior $q_{\phi}$ to be a mixture of $K$ multivariate Gaussians,
$\begin{split}q_{\bm{\phi}}(\mathbf{x})=\sum_{k=1}^{K}w_{k}\mathcal{N}\left({\mathbf{x}};{\bm{\mu}_{k}},{\sigma_{k}^{2}\mathbf{\Sigma}_{\bm{\lambda}}}\right)\end{split}$ (20)
where $\mathbf{\Sigma}_{\bm{\lambda}}\equiv\text{diag}[{{\lambda}^{(1)}}^{2},\ldots,{\lambda^{(D)}}^{2}]$ and $\bm{\lambda}=({\lambda}^{(1)},\ldots,\lambda^{(D)})$ is a vector of parameter scales shared across mixture components; $w_{k}$, $\bm{\mu}_{k}$, and $\sigma_{k}$ are, respectively, the mixture weight, mean, and global scale of the $k$-th component. This choice of variational posterior is both flexible and conducive to enabling (sparse) Bayesian quadrature in the subsequent steps.
@@
$\begin{split}\text{Cov}(\mathcal{I}_{i},\mathcal{I}_{j})&=\int\int\mathcal{N}(\tilde{\mathbf{x}};\mu_{i},\Sigma_{i})\mathcal{N}(\tilde{\mathbf{x}}^{\prime};\mu_{j},\Sigma_{j})\\ &\quad\times\text{Cov}(f(\tilde{\mathbf{x}}),f(\tilde{\mathbf{x}}^{\prime}))\,d\tilde{\mathbf{x}}\,d\tilde{\mathbf{x}}^{\prime}\\ &=\int\int\mathcal{N}(\tilde{\mathbf{x}};\mu_{i},\Sigma_{i})\mathcal{N}(\tilde{\mathbf{x}}^{\prime};\mu_{j},\Sigma_{j})\\ &\quad\times\kappa_{\psi}(\tilde{\mathbf{x}},\tilde{\mathbf{x}}^{\prime})\,d\tilde{\mathbf{x}}\,d\tilde{\mathbf{x}}^{\prime},\end{split}$ (23)
where $\kappa_{\bm{\psi}}$ is the sparse GP posterior covariance function as in Eq. 19. Both the posterior mean and variance of the integral can be obtained in closed form as integrals for the product of Gaussians. Therefore, we can compute the elbo and its standard deviation $\textsc{elbo}_{\textrm{sd}}$ efficiently. Thus, once the sparse GP fitting is done, obtaining a tractable posterior approximation is computationally cheap and robust. Derivations are provided in Supplementary Material A.4.